Circular No. 7605

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ai IN NGC 5278 M. Modjaz and W. D. Li, University of California at Berkeley, on behalf

of LOTOSS (cf. IAUC 7514), report the discovery of an apparent supernova on unfiltered CCD images taken on Mar. 28.4 (mag ∼ 17.6) and 29.4 UT (mag ∼ 17.5) with the 0.8-m Katzman Automatic Imaging Telescope. SN 2001ai is located at α = 13^{h}41^{m}39^{s}.35, δ = +55^{o}40′05′′.1 (equinox 2000.0), which is 2′′.0 west and 9′′.0 south of the nucleus of NGC 5278. An unfiltered image taken by M. Schwartz with the Tenagra III 0.5-m automatic telescope on Mar. 19.4 showed nothing at this position (limiting mag ∼ 19.0). NGC 5278 is one of an interacting connected pair of galaxies (with the center of NGC 5279 ∼ 0′.5 away from that of NGC 5278).

T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of SN 2001ai, obtained by P. Berlind on Mar. 29.25 UT with the F. L. Whipple Observatory 1.5m telescope (+ FAST spectrograph), shows it to be a type-Ic supernova before maximum. The spectrum is similar to that of SN 1994I, ∼ 5 days before V maximum (see Figure 1 of Filippenko et al. 1995, Ap.J. 450, L11), although it is much more blue. This may be due to contamination by stars associated with a very bright underlying H II region. Adopting the NASA/IPAC Extragalactic Database recession velocity of 7541 km/s for the host galaxy, the supernova expansion velocity is ∼ 11 700 km/s for Ca II (rest 395.1 nm).

COMET C/2001 A2 (LINEAR) This comet has apparently undergone a rapid brightening (ephemeris

on IAUC 7600). M. Mattiazzo, Wallaroo, South Australia, notes that the total visual magnitude has brightened by ∼ 2.5 mag in the 24 hr ending Mar. 30.5 UT, with the comet becoming noticeably more condensed in the same period. Visual m_{1} estimates and coma diameters: Mar. 26.82, 10.8, 2′.5 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 28.44, 10.7, 3′.5 (Mattiazzo, 0.20-m reflector); 28.98, 10.9, 3′ (P. M. Raymundo, Salvador, Brazil, 0.25-m reflector); 29.27, 10.9, 2′.0 (M. Linnolt, Honolulu, HI, 0.25-m reflector); 29.46, 10.8, 3′.0 (Mattiazzo); 29.94, 9.5, 3′.5 (A. Amorim, Florianopolis, Brazil, 0.14-m reflector); 30.45, 8.6, 2′.9 (Y. Nagai, Yamanashi, Japan, 0.32-m reflector); 30.40, 8.2, 5′ (S. T. Rae, Whakatane, New Zealand, 10×50 binoculars); 30.52, 8.0, 3′.0 (Mattiazzo, McLaren Vale, S. Australia, 25×100 binoculars); 30.81, 7:, 5′ (K. Černis, Vilnius, Lithuania, no instrument given).

2001 March 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green